How many elements does A=0 have ?


1
Expert's answer
2021-07-12T15:57:30-0400

"P(A)" is the power set of set "A."

Number of elements of "P(A)=2^n," where "n" is the number of elements of the set "A."

Given "A=\empty," then the number of elements of set "A=0."

Number of elements of "P(A)=2^n=2^0=1."

Hence "P(A)" has one element.